40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 16003     Entry time: Wed Apr 7 02:50:49 2021
Author: Koji 
Type: Update 
Category: SUS 
Subject: Flange Inspections 

Basically I went around all the chambers and all the DB25 flanges to check the invac cable configurations. Also took more time to check the coil Rs and Ls.

Exceptions are the TTs. To avoid unexpected misalignment of the TTs, I didn't try to disconnect the TT cables from the flanges.

Upon the disconnection of the SOS cables, the following steps are taken to avoid large impact to the SOSs

  • The alignment biases were saved or recorded.
  • Gradually moved the biases to 0
  • Turned off the watchdogs (thus damping)

After the measurement, IMC was lock and aligned. The two arms were locked and aligned with ASS. And the PRM alignment (when "misalign" was disengaged) was checked with the REFL CCD.
So I believe the SOSs are functioning as before, but if you find anything, please let me know.
 

ELOG V3.1.3-